OpenStudy (abmon98):

Find the equation of the straight line which passes through the point (3,1) and has a gradient of -4. Given that this line is parallel to the tangent to the curve y=k/x^2 at the point where x=2 calculate the value of k

OpenStudy (anonymous):

The gradient cannot be a single number (-4), since it is a vector quantity.

hartnn (hartnn):

gradient = slope = m= -4 x1,y1 = 3,1 so using the slope point form of the equation of line, \(\large y-y_1= m(x-x_1)\) y- 1 = (-4) (x-3) you can simplify this to get your equation of line :)

hartnn (hartnn):

"tangent to the curve y=k/x^2 at the point where x=2" has the slope same as derivative of y at x=2 so calculate dy/dx first, y= k/x^2 = kx^-2 dy/dx = k (-2) x^(-1) now the slope m = dy/dx at x= 2 = k (-2) * 2^(-1) = -k since this tangent is parallel to your line, their slopes are same so, -k = -4 gives you \(\large \color{red}{k=4}\) :)

OpenStudy (abmon98):

When i looked back at my book it appears to be that k=16 and y+4k=13

hartnn (hartnn):

oh, i did my derivative incorrectly. y = kx^-2 so, dy/dx = k (-2) x^-3 - 4 = -2k (2)^-3 4 = 2k/8 k = 32/2 k=16

OpenStudy (abmon98):

Thank you so much for your help :D

hartnn (hartnn):

you're welcome ^_^